Write and solve a system of equations for a given model / word problem:
2. The owner of a small nursery specializes in iris and day lily starter plants. He has been selling irises at a price of $3.00 per dozen and day lilies at a price of $8.00 per dozen. A landscaper wants to buy 15 dozen of a mixture of day lilies and irises, and she is willing to pay $5.00 per dozen for the mixture. How many of each type should the nursery owner include in the mix?

Answer by stanbon(48568)

The owner of a small nursery specializes in iris and day lily starter plants. He has been selling irises at a price of $3.00 per dozen and day lilies at a price of $8.00 per dozen. A landscaper wants to buy 15 dozen of a mixture of day lilies and irises, and she is willing to pay $5.00 per dozen for the mixture. How many of each type should the nursery owner include in the mix?
-----------------------------------------------
Let number of dozen of irises be "x"; Then number of dozen of day lilies is 15-x
------------
Value of the mix is 3x + 8(15-x) = 5*15 dollars
3x + 120 - 8x = 75
-5x = - 45
x = 9 dozen (# of dozen of irises)
15-x = 6 (# of dozen of day lilies)
=====================================
Cheers,
Stan H.